9.2 Molecules and Charge

2. Indicate the total number of valence electrons in the following ions:

NH4+ PO43− NO2 H3O+

3. In the following structures, one atom has a nonzero formal charge. Locate that atom, and indicate the charge.

4. Draw Lewis structures for the following ions. Show formal charges where appropriate.

PH4+ NO3

BrO CO32−

5. Thiocyanate, SCN, can be drawn in two resonance structures, shown below. Calculate the formal charge on each atom of the two forms.
Answer #1

NHU = 5 + 411)-1 = 8 Pou - = 5+ 416) + 3 = 32 NO. = 5+ 216) + 1 = 18 Hoot = 6+311)-1=8 lit) H - N - Z H = N Atom NHUT – FormaPHY Bro- THI H一P-H - 8:0 H CO²- 30: -Ć - 0 suc= ñ o s=c=N-